The area of a rectangular wall of a barn is 72 square feet. It's length is 10 feet longer than twice its width. Find the length and width of the wall of the barn.
width= ___ feet

Answer by stanbon(75887)

The area of a rectangular wall of a barn is 72 square feet. It's length is 10 feet longer than twice its width. Find the length and width of the wall of the barn.
----
Let the width be "w".
Then length = "2w+10"
-----
Equation:
length*width = Area
(2w+10)*w = 72 sq. ft.
---
2w^2 + 10w - 72 = 0
w^2 + 5w - 36 = 0
----
(w+9)(w-4) = 0
width = 4 ft
length = 2w+10 = 18 ft.
==============================
Cheers,
Stan H.
=============


width= ___ feet